How many liters of 60% solution of sulfuric acid should be added to 10 liters of a 30% solution of sulfuric acid to obtain a 50% solution?
Answer by stanbon(75887)

How many liters of 60% solution of sulfuric acid should be added to 10 liters of a 30% solution of sulfuric acid to obtain a 50% solution?
-----
acid + acid = acid
---
0.60x + 0.30*10 = 0.50(x+10)
----
Multiply thru by 100 to get:
60x + 30*10 = 50x + 50*10
---
10x = 20*10
---
x = 20 liters (amt. of 60% solution:
============================
Cheers,
Stan H.
===========================